OpenStudy (anonymous):

One 10.0 Ω resistor is wired in series with two 5.0 Ω resistors in parallel. A 9.0 V battery supplies power to the circuit. What is the total current through the circuit?

Parth (parthkohli):

\[\begin{align}\rm R_{eq} &= 10 + \dfrac{5\cdot 5}{5 + 5} \\ & = 10 + 2.5\\ & = 12.5\end{align}\]\[\rm I = \dfrac{\mathcal E}{R} = \cdots\]